# Improving Visual Place Recognition Based Robot Navigation Through Verification of Localization Estimates

### v1

@@ V Experiment Results
In this section, we present the results and discussion for Experiments 1 and 2. We then discuss an example study for the historical method. We compare metrics as defined in Section IV-A1 and Section IV-B1 between the baseline system (where all matches are retained), the naive thresholds (as diagnostic indicators), our previous SVM approach [44] (which is trained per environment and per VPR technique), and our proposed MLP system (which is trained only per VPR technique).
@@ V-A Experiment 1 Results
Figure 7 shows the variation in goal error for all missions where the robot reported it had arrived at the goal location; Table I details the mean, median and maximum goal errors. Overall for AP-GeM, the proposed MLP verification system reduces the mean goal error by $\approx 73\%$ from the baseline system in completed missions, from 12.51m to 3.42m; compared to the SVM, $N_{P}$ and $N_{R}$ which see $\approx 46\%$, $\approx 69\%$ and $\approx 43\%$ reductions respectively. This trend holds for the NetVLAD and SALAD techniques, which see $\approx 65\%$ and $\approx 68\%$ reductions in mean goal error from the baseline using our proposed system. Furthermore, our MLP network outperforms both naive thresholds, indicating that it has learnt statistical feature input weights that more accurately correlate with VPR integrity than simply using match distances. Figure 7 demonstrates that the verification of VPR is especially beneficial for navigation decisions in adverse conditions, with all VPR feature types showing a significant drop in median error along the logarithmically scaled axis.
@@ V-B Experiment 2 Results
Figure 8: Top-down example of the two improvement modes when utilizing the historical method, between baseline VPR (left) and our proposed system (right) for AP-GeM in Office Adverse. Note that, for readability, the y-scale has been magnified. Black, red, and green correspondences indicate unchanged matches, rejected matches, and corrected matches, respectively, by our proposed verification system.
@@
When including the naive thresholds in the analysis, it initially appears that $N_{P}$ outperforms our proposed system across all metrics in Table II. However, by examining Table III, we see our proposed system consistently achieves the desired balance of performance improvement with data retention (recall), whereas $N_{P}$ discards between ${\approx}75\%$ to ${\approx}53\%$ of the opportunities to localize even when given $1.5m$ of match history to use. Especially considering the threshold must be selected ahead of deployment, there is no guarantee for consistent rejection of out-of-tolerance matches throughout a data set; their large variability is unsuitable for safe navigation systems. Additionally, identifying a performance point for real deployment creates another layer of complexity. An ideal verification system must be adaptive to accurately reject and retain data. As is evident in the present metrics, our proposed system succeeds in removing out-of-tolerance matches regardless of match distance, surpassing all comparisons for precision whilst balancing the retention of matches as reflected in the recall scores.
@@ VI Conclusion
Providing metrics that convey confidence or uncertainty will continue to be a key requirement for trustworthy systems across robotics. This letter presents a real-time integrity-based verification method that can reject VPR errors which may lead to unsafe navigation decisions. We demonstrate two implementations: (a) an instantaneous rejection method in the context of a robot navigating to a goal zone, and (b) a historical method that takes a best, verified match and uses an odometer to extrapolate forwards to a position estimate.
@@
Although our proposed system cannot yet quantitatively estimate error (either lateral or longitudinal), there already exists scenarios where the presented capabilities can be useful. They may provide redundancy, correcting signals, or trigger events such as loop closure in SLAM systems. With our proposed system serving as a demonstrator, further development could see improved heuristics and feature selection for MLP training, training data curation across larger sets, error-binned classes, a weighting scheme, or advanced neural network implementations. For the historical approach, applying weights to history entries may improve performance, which could be either hand-crafted from heuristics or learnt. An example would be promoting recent matches over older matches, which may be especially useful for applications with long histories.

### v2

@@ V Experiment Results
In this section, we present the results and discussion for the two experiments and some additional studies. We compare metrics as defined in Sections IV-A1 and IV-B1 between the baseline system (where all matches are retained), the naive thresholds (as diagnostic indicators), our previous SVM approach [40] (trained per environment and per VPR technique), and our proposed system (trained only per VPR technique).
@@ V-A Experiment 1 Results
Figure 7 shows the variation in goal error for all missions where the robot reported it had arrived at the goal location; Table I details the mean, median and maximum goal errors. Overall for AP-GeM, the proposed MLP verification system reduces the mean goal error by $\approx 73\%$ from the baseline system in completed missions, from 12.51m to 3.42m; compared to $N_{P}$, $N_{R}$, and the SVM, which see $\approx 46\%$, $\approx 69\%$ and $\approx 43\%$ reductions respectively. This trend holds for the NetVLAD and SALAD techniques, which see $\approx 65\%$ and $\approx 68\%$ reductions in mean goal error from the baseline using our proposed system. Furthermore, our MLP network outperforms both naive thresholds, indicating that it has learnt statistical feature input weights that more accurately correlate with VPR integrity than simply using match distances. Figure 7 demonstrates that the verification of VPR is especially beneficial for navigation decisions in adverse conditions, with all VPR feature types showing a significant drop in median error along the logarithmically scaled axis.
@@ V-B Experiment 2 Results
Figure 8: E.g. of the two improvement modes when utilizing the HoQ method, between baseline VPR (left) and our proposed system (right) for AP-GeM in Office Adverse. Note: For readability, the y-scale has been magnified. Black, red, and green correspondences denote unchanged matches, rejected matches, and corrected matches, respectively, by our proposed verification system.
@@
When including the naive thresholds in the analysis, $N_{P}$ initially appears to outperform our proposed system across all metrics in Table II. However, by examining Table III, we see our proposed system consistently achieves the desired balance of performance improvement with data retention (recall), whereas $N_{P}$ discards between ${\approx}75\%$ to ${\approx}53\%$ of the opportunities to localize even with $1.5m$ of match history. Especially considering the threshold must be selected ahead of deployment, there is no guarantee for consistent rejection of out-of-tolerance matches throughout a data set; their large variability is unsuitable for safe navigation systems. Additionally, identifying a performance point for real deployment creates another layer of complexity. An ideal verification system must be adaptive to accurately reject and retain data. As is evident in the metrics, our proposed system succeeds in removing out-of-tolerance matches regardless of match distance, surpassing all comparisons for precision whilst balancing the retention of matches as shown in the recall scores.
@@ V-C Example Study: Experiment 2 Training Generalizability
V-D Stand Alone VPR Performance
Table V provides precision, recall@1, and accuracy values, as calculated in [3], for VPR performance using NetVLAD, aggregated across the QCR datasets, and for two completely unseen datasets: Nordland [43] (winter reference and summer query sets) and Oxford RobotCar [44] (sunny reference and overcast query sets). Nordland queries are sampled 1:10 with reference set from [43], Oxford RobotCar sets are sampled every 1m. Here, we use the same models for our single-query method and the SVM from the above experiments, trained on the QCR datasets. The baseline performance is equal in all metrics because it accepts all VPR matches. Values surpassing the baseline are in bold, otherwise the best verification method is underlined. As seen, our verification system, in general, has the highest precision and accuracy, with varied reductions in recall. This indicates the model has some generalizability to data from a totally different distribution to the training data.
TABLE V: Precision, recall@1, and accuracy for VPR performance across datasets. The baseline comparison ‘accepts’ all VPR position estimates. The models trained for the above experiments on the QCR training sets are used across all datasets. We compute precision and recall@1 for VPR as in [3].
Net- QCR Aggregate Nordland [43] RobotCar [44]
VLAD VPR SVM Ours VPR SVM Ours VPR SVM Ours
Prec. 75.4% 77.8% 87.2% 8.6% 7.0% 13.6% 77.0% 71.4% 83.1%
Rec@1 75.4% 24.3% 49.7% 8.6% 6.8% 4.02% 77.0% 53.4% 67.9%
Acc. 75.4% 41.9% 69.0% 8.6% 8.3% 69.8% 77.0% 54.9% 77.1%
@@ VI Conclusion
Providing metrics that convey confidence or uncertainty will continue to be a key requirement for trustworthy systems across robotics. This letter presents a real-time integrity-based verification method that can reject VPR errors which may lead to unsafe navigation decisions. We demonstrate two implementations: (a) a single-query rejection method in the context of a robot navigating to a goal zone, and (b) a history-of-queries method that takes a best, verified match and uses an odometer to extrapolate forwards to a position estimate.
@@
Although our proposed system cannot yet quantitatively estimate error (either lateral or longitudinal), there already exists scenarios where the presented capabilities can be useful. They may provide redundancy, correcting signals, or trigger events such as loop closure in SLAM systems. With our proposed system serving as a demonstrator, further development could see improved heuristics and feature selection for MLP training, training data curation across larger sets, error-binned classes, a weighting scheme, or advanced neural network implementations. For the HoQ method, applying weights to history entries may improve performance, which could be either hand-crafted from heuristics or learnt. An example would be promoting recent matches over older matches, which may be especially useful for applications with long histories.